OpenStudy (anonymous):

Rewrite the following expression with positive exponents. (28xy)^-9/10

OpenStudy (anonymous):

one step, recall that \[b^{-n}=\frac{1}{b^n}\]

OpenStudy (anonymous):

you get \[\frac{1}{(28xy)^{\frac{9}{10}}}\]